Zeus Shoe Limited is a manufacturer of ladies shoes and uses a standard costing system. The standard manufacturing overhead cost per shoe are based on direct labour hours and are as follows:

Variable Overhead (5 hours @ $24 per hour)                                     $120

Fixed Overhead (5 hours @ $38 per hour)                                         $180

                        TOTAL OVERHEAD                                               $310

(Based on capacity of 300,000 direct labour hours per month)

The following information is available for the month of October.

56,000 shoes were produced, although 60,000 shoes were budgeted

275,000 direct labour hours were worked at a total cost of $7,650,000

Variable overhead cost were $7,020,000

Fixed cost overhead were $1,250,000

REQUIRED

Calculate the

Variable overhead spending variance                                     5 Marks

Variable overhead efficiency variance                                    5 Marks

Fixed overhead budget variance                                             5 Marks

Fixed overhead volume variance                                            5 Marks
Answer #1

1. Variable overhead spending variance = Actual spending on variable overheads - Budgeted spending on variable overheads

= 7020000 - (275000*24)

= 7020000 - 6600000

= $420000 Unfavorable

2. Variable overhead efficiency variance = Standard variable overhead rate x (Actual hours - Standard hours)

= 24*[275000-(56000*5)]

= 24*(-5000)

= $ 120000 Favorable

3. Fixed overhead budget variance = Actual Fixed Overhead - Budgeted Fixed Overhead

= 1250000 - (60000*5*38)

= 1250000 - 11400000

= $10150000 Favorable

4. Fixed overhead volume variance = Budgeted Fixed overhead - Standard Overhead

= 60000*190 - 56000*38*5

= 11400000 - 10640000

= $760000 Favorable